OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle.

OpenStudy (anonymous):

Let x and y be the two complementary angles since they are complementary, we know they add to 90 degrees, so x+y = 90

OpenStudy (anonymous):

we are told that "The sum of the measures of two complementary angles exceeds the difference of their measures by 72", so x+y = x-y + 72 y = -y + 72 y + y = 72 2y = 72 y = ???

OpenStudy (anonymous):

Complementary angles have measures that add up to 90. Let one angle have measure x. The other angle has measure 90 - x Sum of measures: x + (90 - x) Difference of their measures: x - (90 - x) Sum of measures exceeds diff of measures by 72: x + (90 + x) = x - (90 - x) + 72 Solve for x

OpenStudy (anonymous):

yo dawg dat s hit is f ucking hard i rather just guess or u could give me da awnser